b) In the step index fiber the relative refractive index difference is=4% and n2 = 1.5 Find a) n1 b)Critical propagation angle c)N.A d)Maximum angle of incidence for guidance
1
Expert's answer
2020-05-25T11:18:39-0400

Find "n_1" from relative refractive index difference "\Delta":


"\Delta=\frac{n_1^2-n_2^2}{2n_1^2},\\
n_1=1.56."

Calculate the numerical aperture:


"NA=\sqrt{n_1^2-n_2^2}=0.428."

Calculate critical propagation angle:


"\theta_c=\text{sin}^{-1}\frac{n_1}{n_2}=74.1^\circ."

Maximum angle of incidence (or acceptance angle):


"\theta_{acc}=\text{sin}^{-1}\bigg(\frac{1}{n_0}\sqrt{n_1^2-n_2^2}\bigg)=25.4^\circ."

In this equation, "n_0" is the refractive index of a medium around the fibre. It is close to unity for air.